OpenStudy (anonymous):

How do you find the differential of the integral cos^3(x) dx from -9ts to -5t+6s?

OpenStudy (anonymous):

\[\int\limits_{-9ts}^{-5t+6s}\cos ^{3}(x) dx\]

OpenStudy (amistre64):

ummmm......... i dont really know...

OpenStudy (amistre64):

i spose its possible, just use those"equations" as the F(x) stuff in the end

OpenStudy (amistre64):

you might want to go about "reducing" the integrand to an addition of sin and cos....maybe

OpenStudy (amistre64):

cos cos^2 cos(1-sin^2) cos(x)(1 - (1-cos(2x)/2)) ... along those lines eh?

OpenStudy (dumbcow):

yeah thats what i was going to say

OpenStudy (amistre64):

well say it again then :)

OpenStudy (dumbcow):

integral of sin^2 is known as -x/2 + sincos/2 so you could use that

OpenStudy (anonymous):

Well I have notes and an example problem, but it doesnt really make sense. He has the differential of the integral e^-x^2 from a to b. He then takes partials. Fa = -e^a^2 and Fb = e^b^2. Then apparently the differential is -e^a^2 * da + e^b^2 * db. No sense.

OpenStudy (anonymous):

Don't worry about it, I just wondered if anyone could make sense out of it, but obviously my teacher is drunk. He wrote his own "textbook" in Mathematica and he didn't do a very good job.

OpenStudy (dumbcow):

are you still talking about integral of cos^3? anyway yeah there are no elementary functions to describe integral of e^(-x^2).